OpenStudy (anonymous):

The number of men in a tour group is 15 times the number of women. If there are 48 people in the group, how many are men?

OpenStudy (anonymous):

45 30 15 10

OpenStudy (anonymous):

well let's imagine that the number of men = x and the number of women = y we have x = (15) y and x + y = 48

OpenStudy (anonymous):

ok then?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so u have this system: x = 15y x + y = 48 you have x = 15y so you can say 15y + y = 48 i just plugin x = 15y in the second equation.

OpenStudy (anonymous):

can you the rest?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ok lets start with 5 ok so try (5*15)+5

OpenStudy (anonymous):

@usitdown just you need to solve for y: 15y + y = 48

OpenStudy (anonymous):

does that work?

OpenStudy (anonymous):

@usitdown you there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im confussed on what pfeh is saying

OpenStudy (anonymous):

ya he tryed to make it more complicated than it actually is or needs to be

OpenStudy (anonymous):

though his way works very well to and will work for most harder one of this type too

OpenStudy (anonymous):

let's start from the beginning,we imagined that the number of men = x and the number of women = y so we had x = 15y and in the question it was written that the number of people is 48 so x + y = 48

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

do you need my help or has @PFEH.1999 go it?

OpenStudy (anonymous):

and we know that x = 15y,besides this we know that x + y = 48 so we can easily write : (15y) + y = 48

OpenStudy (anonymous):

Any clue what y equals

OpenStudy (anonymous):

try one 1,2,3,4,5,6,7,8,9

OpenStudy (anonymous):

The 4 answer choices are at the top

OpenStudy (anonymous):

But they seem all to big

OpenStudy (anonymous):

yes they're incorrect!

OpenStudy (anonymous):

no there not

OpenStudy (anonymous):

No they arent because >>>>>>THOSE ARE THE ANSWER CHOICES<<<<<<<

OpenStudy (anonymous):

one of them is correct

OpenStudy (anonymous):

ok dont listen to @PFEH.1999

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

oh no! i made a mistake! you want the number of men! i thought you want the number of women

OpenStudy (anonymous):

Yea @helper99

OpenStudy (anonymous):

ok so what 15*1+2 ?

OpenStudy (anonymous):

15*1+2=17

OpenStudy (anonymous):

hurry!!

OpenStudy (anonymous):

so that does not work so lets try 15*2+2

OpenStudy (anonymous):

=32

OpenStudy (anonymous):

15*3+3=47

OpenStudy (anonymous):

48*

OpenStudy (anonymous):

48-3=45 which is the answer